b) The wait time customers experience when calling into a computer help hotline is distributed uniformly between 0 and 15 minutes. Find the expected wait time a user will experience. (Round your answer to one decimal place)

c) Assume that the heights of students at a high school are normally distributed, with mean 65 inches and standard deviation 2.5 inches. If 5 students are randomly selected, what is the probability that fewer than 2 are taller than 65 inches? (Round your answer to four decimal places)
Answer #1

a)

here for uniform distribution parameter a =0 and b=15
mean wait time μ = (a+b)/2 = 7.5

b)

here as 65 is at mean, and being symmetric distribution: probability a student is taller than 65 inches =0.5

probability that fewer than 2 are taller than 65 inches =P(X<2) =P(X=0)+P(X=1)=(5C0)*(0.5)0(0.5)5+(5C1)*(0.5)1(0.5)4

=0.1875